# Optimized Multi-Token Joint Decoding with Auxiliary Model for LLM Inference

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable success across diverse tasks, yet their inference processes are hindered by substantial time and energy demands due to single-token generation at each decoding step. While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution, thereby enhancing speed without improving output quality. In contrast, our work simultaneously enhances inference speed and improves the output effectiveness. We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and enhancing task performance. However, MTJD suffers from the high cost of sampling from the joint distribution of multiple tokens. Inspired by speculative decoding, we introduce multi-token assisted decoding (MTAD), a novel framework designed to accelerate MTJD. MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also improves the decoding efficiency over conventional speculative decoding. Theoretically, we demonstrate that MTAD closely approximates exact MTJD with bounded error. Empirical evaluations using Llama-2 and OPT models ranging from 13B to 70B parameters across various tasks reveal that MTAD reduces perplexity by 21.2% and improves downstream performance compared to standard single-token sampling. Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than conventional speculative decoding methods. These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more sustainable and high-performance deployment of LLMs.
@@ 1 Introduction
To tackle this challenge, researchers have delved into non-autoregressive decoding approaches. Early methods (Ghazvininejad et al., 2019; Gu et al., 2017; Guo et al., 2020) aimed at reducing inference latency by concurrently generating multiple tokens. But these methods usually require task-dependent techniques and information to match the performance of autoregressive decoding (Kim et al., 2023; Xiao et al., 2023). More recently, speculative decoding has emerged (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023), exploiting the observation that most of the small model’s prediction aligns well with that of a large model. It leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model. As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding achieves an overall speed-up of $1$-2×2$\times$2 ×. Despite gains in speed, these methods still generate each token based on its single-token probability. Consequently, it does not enhance the effectiveness of the generated sequences.
In this work, we aim to go beyond the conventional trade-off between efficiency and effectiveness by introducing multi-token joint decoding (MTJD). Unlike traditional approaches, MTJD produces multiple tokens from their joint distribution at each decoding step. Theoretically, we show this joint generation can lead to lower perplexity and hence improved task performance. However, directly sampling from the joint distribution of multiple tokens poses significant computational challenges, rendering MTJD impractical.
Inspired by speculative decoding, we propose multi-token assisted decoding (MTAD), a novel framework designed to approximate and accelerate MTJD. MTAD employs a smaller auxiliary model to estimate the joint distribution of a larger model, significantly reducing computational demands. To ensure the accuracy of this approximation, MTAD incorporates a verification mechanism that not only guarantees the accuracy of the draft tokens but also enhances efficiency beyond conventional speculative decoding by maximizing the number of accepted tokens per iteration. We provide both theoretical and empirical analyses to demonstrate that MTAD improves perplexity and downstream performance. Meanwhile, it achieves significant reductions in energy and time usage compared to existing decoding strategies.
@@
    We analyze the energy consumption of LLM inference. To our knowledge, we are the first to give theoretical and empirical evidence that, despite that MTAD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they reduce the overall energy consumption by reducing the overhead induced by accessing GPU global memory.
@@
    We conducted comprehensive evaluations with Llama-2 and OPT models (ranging from 13B to 70B parameters) across various tasks, demonstrating that MTAD enhances perplexity by 21.2% and improves downstream effectiveness compared to standard single-token sampling, while also achieving a 1.42$\times$ speed-up and reducing energy consumption by 1.54$\times$ compared to conventional speculative decoding methods.
These advancements position MTAD as a robust solution for making multi-token joint decoding both effective and efficient, thereby facilitating more sustainable and high-performance deployment of large-scale language models. Our code is publicly available [^1] https://anonymous.4open.science/r/LLMSpeculativeSampling-EE52.
@@ 2.1 Decodings of LLMs
Let $p$ denote the distribution defined by LLM model $M_{p}$. Given an input context $input$, a decoding algorithm generates a sequence of $N$ tokens whose likelihood is denoted as $p(x_{1:N}|input)$. The likelihood of the sequence is directly linked to perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens. Based on autoregressive decomposition $p(x_{1:N}|input)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},input)$ [^2] In the paper, we omit $input$ when there is no ambiguity., the perplexity is defined as:
@@
Perplexity serves as a direct metric for assessing the effectiveness of a decoding algorithm. In practice, when a model is well-trained, lower perplexity often correlates with improved downstream performance. For example, beam sampling aims to return output with lower perplexity and is proven to have better downstream performance in general (Shi et al., 2024).
To further demonstrate the relationship between perplexity and downstream performance, we evaluate GPT-3.5-turbo on the spider (Yu et al., 2018) dataset. Using a temperature of 2, the model generated 10 outputs for each input. We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5-th lowest) perplexity. As shown in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
@@
Multinomial sampling, also known as standarized sampling or single-token sampling, samples the next token $x_{t}$ based on $\mathcal{T}\circ p(\cdot|x_{1:t-1},input)$, where $\mathcal{T}$ is a warping operation applied to enhance the high probability region. Some common warping operations include top-k warping, which limits the selection to the top k tokens, and top-p warping, where tokens are sampled from the smallest possible subset of the vocabulary whose cumulative probability mass exceeds a specified threshold. The deterministic version of multinomial sampling (i.e., greedy decoding) is a special case when $k=1$.
@@
Beam sampling aims to improve output perplexity over multinomial sampling. For each position $t$ ($1\leq t\leq N$), it maintains $W>1$ candidate sequences, which are also called beams. Assume we have already kept the $W$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(W)}\}$ at position $t-1$, $W$ sequences with length $t$ are then sampled from $\mathcal{T}\circ p_{beam}$, where $p_{beam}:\mathcal{I}_{t-1}\times V\rightarrow[0,1]$ is the beam sampling probability:
$p_{beam}(x_{1:t-1}^{(i)},x_{t})=\frac{p(x_{1:t-1}^{(i)},x_{t}|input)}{\sum_{x_{1:t-1}^{(j)},x^{\prime}_{t}\in\mathcal{I}_{t-1}\times V}p(x_{1:t-1}^{(j)},x_{t}^{\prime}|input)}$ (2)
@@ 2.2 Vanilla Speculative Decoding
    Let $input$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ using multinomial sampling based on $\tilde{q}(x_{t}|x_{1:t-1},input))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
@@
    Repeat step 1-3 until reaching the stopping criteria, e.g., reaching the length limit..
Because the large model verifies $\gamma$ tokens in parallel with one run, the time cost is smaller than calling it $\gamma$ times. Meanwhile, although the small model still runs in an autoregressive way, its inference speed is much faster than the large model. As a result, speculative decoding achieves a speedup of 1–2× compared to multinomial sampling while maintaining an identical sampling distribution.
@@ 3 Methodology
As discussed in Section 2, the goal of this work is to design an algorithm that yields lower perplexity and better efficiency than multinomial sampling and vanilla speculative decoding. In this section, we first introduce multi-token joint decoding (MTJD), which generates multiple tokens based on their joint likelihood. We prove it can yield lower perplexity. Then we introduce multi-token assisted decoding (MTAD), which approximates and accelerates MTJD by exploiting an auxiliary model.
@@ 3.1 Multi-Token Joint Decoding
We first introduce a new decoding algorithm to improve multinomial sampling in terms of perplexity.
@@
Multi-Token Joint Decoding. Let $M_{p}$ be the large target model with distribution $p$. Different from single-token multinomial sampling, multi-token joint decoding (MTJD) generates the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$, where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
@@
Theorem 3.2 shows the limit of perplexity of MTJD when $N$ approaches infinity. The proofs are included in the Appendix A.
@@
Empirical evidence supports our claim. We fine-tune both a Llama and an OPT model on the ChatGPT-Prompts dataset and evaluate the output perplexity and Rouge-L scores with example outputs. Figure 1 shows the output perplexity and Rouge-L scores of MTJD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$. Notice that setting $K=1$ is equivalent to multinomial sampling. We use beam sampling to approximate the $\arg\max$ sampling from the joint distribution $p(x_{t+1:t+K}|x+{1:t},input)$. We can see that the perplexity keeps dropping when $K$ increases. It confirms our claim that increasing $\gamma_{i}$ will increase the output perplexity. Moreover, the Rouge-L score also improves with $K$, supporting our claim that better perplexity reflects enhanced performance in downstream tasks.
@@ 3.2 Multi-Token Assisted Decoding
Unfortunately, the computation cost of MTJD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$. Inspired by speculative decoding and the facts that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” (Leviathan et al., 2023; Kim et al., 2023), we propose multi-token assisted decoding (MTAD), which exploits a small auxiliary model $M_{q}$ to accelerate MTJD approximately. The core idea is to (1) use the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$ output by $M_{q}$ to approximate $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^3] It is also valid to approximate $\tilde{p}$ with $\tilde{q}$. Without loss of generality, we consider non-warped distribution in the illustration of MTAD. and generate $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) use the large model to validate draft tokens in parallel and accept the longest draft prefix sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration. However, it is still infeasible to directly generate draft tokens from the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$. So we propose to further approximate this process with beam sampling, which is an effective and efficient algorithm to generate sequences with high likelihood. In this way, MTAD reduces the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does. Algorithm 1 in the Appendix illustrates the pseudocode of MTAD algorithm.
@@
Figure 2 illustrates the verification process of MTAD. Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam sampling with the auxiliary model. Since beam sampling is a widely recognized algorithm to generate sequences with high overall likelihood (Leblond et al., 2021), it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large. Also, since beam sampling works in an autoregressive way, we can also assume that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large. To approximate MTJD, for each step $i$, MTAD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$. So MTAD first computes the joint likelihood $p(x_{t+1:t+j}|x_{1:t})$ for $j=1,\ldots,\gamma$. Then for each prefix sub-sequence $x_{t+1:t+j}$, it passes verification if and only if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, where $\tau\in[0,1)$ is a pre-defined threshold. Notice that if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, we have $\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})}>\tau$, which means $\frac{q(x_{t+1:t+j}|x_{1:t})-p(x_{t+1:t+j}|x_{1:t})}{p(x_{t+1:t+j}|x_{1:t})}<\frac{1}{\tau}-1$. Therefore, our acceptance policy guarantees that when $q(x_{t+1:t+j}|x_{1:t})>p(x_{t+1:t+j}|x_{1:t})$, the relative error is bounded. And if $q(x_{t+1:t+j}|x_{1:t})\leq p(x_{t+1:t+j}|x_{1:t})$, it means the sub-sequence has higher likelihood in the large model, then it is reasonable to accept it. After verifying all the sub-sequences, MTAD accepts the longest prefix sub-sequence that passes verification.
The verification step of MTAD ensures that the accepted tokens have a high joint likelihood with the large model. We have shown that selecting multiple tokens based on their joint likelihood lead to better output perplexity. Thus, MTAD is more effective than multinomial sampling and vanilla speculative decoding. Furthermore, since MTAD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high. So at each iteration, MTAD can accept more draft tokens than vanilla speculative decoding, which results in better efficiency.
Next, we theoretically analyze the approximation error of MTAD. Lemma 3.4 shows the upper bound of MTAD’s perplexity. And Theorem 3.5 shows the upper bound of the ratio between the perplexity of approximate MTAD and exact MTJD. The proofs are given in Appendix A.
@@
$\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))\geq(1-\epsilon)\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$ (5)
@@
With the two assumption above, when $N\rightarrow\infty$ we have
@@
With Theorem 3.6, we observe that when $q\rightarrow p$, we have $U\rightarrow 0$ and $\bar{\gamma}\rightarrow\infty$. Meanwhile, when $\epsilon\rightarrow 0$, meaning the beam width for the auxiliary model is large enough, the ratio bound in Theorem 3.5 converges to 1, It implies that MTAD converges to MTJD under these limiting conditions.
Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to generate multiple draft sequences and applying tree-based attention (Miao et al., 2023) for simultaneous verification. However, our preliminary experiment results suggest that since MTAD already selects the longest accepted prefix sub-sequence, the advantage of generating multiple draft tokens is less significant. Moreover, this approach increases the memory cost during inference and may affect the error bounds derived above. Therefore, we leave a more detailed exploration of this extension as future work.

### v3

@@ Abstract
Large language models (LLMs) have achieved remarkable success across diverse tasks, but, due to single-token generation at each decoding step, their inference processes are hindered by substantial time and energy demands. While previous methods such as speculative decoding mitigate these inefficiencies by producing multiple tokens per step, each token is still generated by its single-token distribution. Although this enhances the speed, it does not improve the output quality. In contrast, our work simultaneously boosts inference speed and improves the output effectiveness. We consider multi-token joint decoding (MTJD), which generates multiple tokens from their joint distribution at each iteration, theoretically reducing perplexity and raising task performance. However, MTJD suffers from the high cost of sampling from the joint distribution of multiple tokens. Inspired by speculative decoding, we introduce multi-token assisted decoding (MTAD), a novel framework designed to accelerate MTJD. MTAD leverages a smaller auxiliary model to approximate the joint distribution of a larger model, incorporating a verification mechanism that not only ensures the accuracy of this approximation, but also increases the decoding efficiency over conventional speculative decoding. Theoretically, we demonstrate that MTAD closely approximates exact MTJD with a bounded error. Empirical evaluations across various tasks reveal that MTAD improves downstream performance by 25% compared to standard single-token sampling. Furthermore, MTAD achieves a 1.42$\times$ speed-up and consumes 1.54$\times$ less energy than vanilla speculative decoding methods. These results highlight MTAD’s ability to make multi-token joint decoding both effective and efficient, promoting more productive and high-performance deployment of LLMs. [^1] We release our code at https://github.com/ZongyueQin/MTAD
@@ 1 Introduction
To tackle this challenge, researchers have delved into non-autoregressive decoding approaches. Early methods (Ghazvininejad et al., 2019; Gu et al., 2017; Guo et al., 2020) aimed at reducing inference latency by concurrently generating multiple tokens. But these methods usually require task-dependent techniques and information to match the performance of autoregressive decoding (Kim et al., 2023; Xiao et al., 2023). More recently, speculative decoding has emerged (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023). This method exploits the observation that most of the small model’s prediction aligns well with that of a large model. It leverages a smaller auxiliary model to draft a few future tokens autoregressively, which are subsequently validated in parallel by the larger model. As the smaller model operates significantly faster and parallel token verification incurs a similar time cost as generating a single token, speculative decoding attains an overall speed-up of $1$-2×2$\times$2 ×. Despite gains in speed, these methods still generate each token based on its single-token probability. Consequently, it does not enhance the effectiveness of the generated sequences.
In this work, we first go beyond the conventional trade-off between efficiency and effectiveness and explore multi-token joint decoding (MTJD). Unlike traditional approaches, MTJD produces multiple tokens from their joint distribution at each decoding step. Theoretically, we show this joint generation can lead to lower perplexity and hence improved task performance. However, directly sampling from the joint distribution of multiple tokens poses significant computational challenges, rendering MTJD impractical.
Inspired by speculative decoding, we propose multi-token assisted decoding (MTAD), a novel framework designed to approximate and accelerate MTJD. MTAD employs a smaller auxiliary model to estimate the joint distribution of a larger model, significantly reducing computational demands. To ensure the accuracy of this approximation, MTAD incorporates a verification mechanism that not only guarantees the accuracy of the draft tokens but also enhances efficiency beyond conventional speculative decoding by maximizing the number of accepted tokens per iteration. We provide both theoretical and empirical analyses to demonstrate that MTAD boosts perplexity and downstream performance. Meanwhile, it significantly reduces the energy and time usage compared to existing decoding strategies.
@@
    We analyze the energy consumption of LLM inference. To our knowledge, we are the first to give quantified and empirical evidence that, despite the fact that MTAD and other speculative decoding algorithms increase the number of FLOPs needed during LLM inference, they actually use less energy with fewer accesses to the GPU global memory.
@@
    We conducted comprehensive evaluations across various tasks, demonstrating that MTAD improves downstream performance by 25% compared to standard single-token sampling, while, at the same time, realizing a 1.42$\times$ speed-up and paring energy consumption by 1.54$\times$ compared to vanilla speculative decoding methods.
These advancements position MTAD as a robust solution for making multi-token joint decoding both effective and efficient, thereby facilitating more sustainable and high-performance deployment of large-scale language models.
@@ 2 Preliminaries
In this section, we discuss preliminaries relevant to contextualizing our paper.
@@ 2.1 Decodings of LLMs
Let $p$ denote the distribution defined by LLM model $M_{p}$. Given an input context $input$, a decoding algorithm generates a sequence of $N$ tokens whose likelihood is designated as $p(x_{1:N}|input)$. The likelihood of the sequence is directly linked to the perplexity of the sequence, which is the exponentiated average negative log-likelihood of all tokens. Based on autoregressive decomposition $p(x_{1:N}|input)=\prod_{t=1}^{N}p(x_{t}|x_{1:t-1},input)$ [^2] In the paper, we omit $input$ when there is no ambiguity., the perplexity is defined as:
@@
Perplexity serves as a direct metric for assessing the effectiveness of a decoding algorithm. In practice, when a model is well-trained, lower perplexity often correlates with improved downstream performance. For example, beam sampling (explained below) aims to return output with lower perplexity and is empirically proven to have better downstream performance in general (Shi et al., 2024).
To further demonstrate the relationship between perplexity and downstream performance, we evaluate GPT-3.5-turbo on the spider (Yu et al., 2018) dataset. Employing a temperature of 2, the model generated 10 outputs for each input. We measured the average perplexities and execution accuracies for the outputs with the highest, lowest, and median (the 5th lowest) perplexity. As seen in Table 1, lower perplexity correlates with improved downstream performance, even in one of today’s largest models.
@@
Multinomial sampling, also known as standarized sampling or single-token sampling, samples the next token $x_{t}$ based on $\mathcal{T}\circ p(\cdot|x_{1:t-1},input)$, where $\mathcal{T}$ is a warping operation applied to enhance the high probability region. Some common warping operations include top-k warping. This limits the selection to the top k tokens, and top-p warping, where tokens are sampled from the smallest possible subset of the vocabulary whose cumulative probability mass exceeds a specified threshold. The deterministic version of multinomial sampling is a special case with $k=1$, also called greedy decoding.
@@
Beam sampling is intended to decrease output perplexity over multinomial sampling. For each position $t$ ($1\leq t\leq N$), it maintains $W>1$ candidate sequences, which are also called beams. Assume we have already kept the $W$ sequences $\mathcal{I}_{t-1}=\{x_{1:t-1}^{(1)},\ldots,x_{1:t-1}^{(W)}\}$ at position $t-1$. $W$ sequences with length $t$ are then sampled from $\mathcal{T}\circ p_{beam}$, where $p_{beam}:\mathcal{I}_{t-1}\times V\rightarrow[0,1]$ is the beam sampling probability:
$p_{beam}(x_{1:t-1}^{(i)},x_{t})=\frac{p(x_{1:t-1}^{(i)},x_{t}|input)}{\sum_{1\leq j\leq W,x^{\prime}_{t}\in V}p(x_{1:t-1}^{(j)},x_{t}^{\prime}|input)}$ (2)
@@ 2.2 Vanilla Speculative Decoding
    Let $input$ be the input context, the small model samples $\gamma$ draft tokens $x_{1},\ldots,x_{\gamma}$ with multinomial sampling based on $\tilde{q}(x_{t}|x_{1:t-1},input))$ for $t=1,\ldots,\gamma$, where $\tilde{q}=\mathcal{T}\circ q$ and $q$ is the small model’s output distribution.
@@
    Repeat step 1-3 until reaching the stopping criteria, e.g., reaching the length limit.
Because the large model verifies $\gamma$ tokens in parallel with one run, the time cost is smaller than calling it $\gamma$ times. Moreover, the global memory access is also pared, which saves energy consumption, as we shall illustrate in Section 4. Meanwhile, although the small model still runs in an autoregressive way, its inference speed is more efficient than the large model. As a result, speculative decoding maintains an identical sampling distribution while realizing a speedup of 1–2$\times$ compared to multinomial sampling and using less energy.
@@ 3 Methodology
As discussed in Section 2, the goal of this work is to design an algorithm that yields lower perplexity and better efficiency than multinomial sampling and vanilla speculative decoding. In this section, we first introduce multi-token joint decoding (MTJD). This generates multiple tokens based on their joint likelihood. We prove it can yield lower perplexity. Then we present multi-token assisted decoding (MTAD), which approximates and accelerates MTJD by exploiting an auxiliary model.
@@ 3.1 Multi-Token Joint Decoding
We first explain a new decoding algorithm to improve multinomial sampling in terms of perplexity.
@@
Multi-Token Joint Decoding. Let $M_{p}$ be the large target model with distribution $p$. Different from single-token multinomial sampling, multi-token joint decoding (MTJD) produces the next $\gamma_{i}$ tokens at step $i$ based on their joint conditional probability $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$, where $\gamma_{i}$ is an integer no less than 1 and $t=\sum_{i^{\prime}=1}^{i-1}{\gamma_{i^{\prime}}}$, i.e., the total tokens generated in the previous $i-1$ steps.
@@
Theorem 3.2 demonstrates the limit of perplexity of MTJD when $N$ approaches infinity. The proofs are included in the Appendix A.
@@
Empirical evidence supports our claim. We fine-tune both a Llama and an OPT model on the ChatGPT-Prompts dataset and evaluate the output perplexity and Rouge-L scores with example outputs. Figure 1 depicts the output perplexity and Rouge-L scores of MTJD with $\gamma_{i}$ set to a constant $K$, where $K=1,\ldots,5$. Notice that setting $K=1$ is equivalent to multinomial sampling. We use beam sampling to approximate the $\arg\max$ sampling from the joint distribution $p(x_{t+1:t+K}|x+{1:t},input)$. We can see that the perplexity keeps dropping when $K$ increases. It confirms our claim that increasing $\gamma_{i}$ will increase the output perplexity. Moreover, the Rouge-L score also improves with $K$, supporting our claim that better perplexity reflects enhanced performance in downstream tasks.
@@ 3.2 Multi-Token Assisted Decoding
Unfortunately, the computation cost of MTJD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$. Inspired by speculative decoding and the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” (Leviathan et al., 2023; Kim et al., 2023), we propose multi-token assisted decoding (MTAD), which exploits a small auxiliary model $M_{q}$ to accelerate MTJD approximately. The core idea is to (1) use the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$ output by $M_{q}$ to approximate $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^3] It is also valid to approximate $\tilde{p}$ with $\tilde{q}$. Without loss of generality, we consider non-warped distribution in the illustration of MTAD. and produce $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) utilize the large model to validate draft tokens in parallel and accept the longest draft prefix sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration. However, it is still infeasible to directly generate draft tokens from the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$. So we propose to further approximate this process with beam sampling, which is an effective and efficient algorithm to generate sequences with high likelihood. In this way, MTAD decreases the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does. Algorithm 1 in the Appendix illustrates the pseudocode of MTAD algorithm.
@@
Figure 2 displays the verification process of MTAD. Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam sampling with the auxiliary model. Since beam sampling is a widely recognized algorithm to produce sequences with high overall likelihood (Leblond et al., 2021), it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large. Also, since beam sampling works in an autoregressive way, we can also infer that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large. To approximate MTJD, for each step $i$, MTAD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$. So MTAD first computes the joint likelihood $p(x_{t+1:t+j}|x_{1:t})$ for $j=1,\ldots,\gamma$. Then for each prefix sub-sequence $x_{t+1:t+j}$, it passes verification if and only if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, where $\tau\in[0,1)$ is a pre-defined threshold. Notice that if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, we have $\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})}>\tau$, which means $\frac{q(x_{t+1:t+j}|x_{1:t})-p(x_{t+1:t+j}|x_{1:t})}{p(x_{t+1:t+j}|x_{1:t})}<\frac{1}{\tau}-1$. Therefore, our acceptance policy guarantees that when $q(x_{t+1:t+j}|x_{1:t})>p(x_{t+1:t+j}|x_{1:t})$, the relative error is bounded. And if $q(x_{t+1:t+j}|x_{1:t})\leq p(x_{t+1:t+j}|x_{1:t})$, it means the sub-sequence has higher likelihood in the large model, then it is reasonable to accept it. After verifying all the sub-sequences, MTAD accepts the longest prefix sub-sequence that passes verification.
The verification step of MTAD ensures that the accepted tokens have a high joint likelihood with the large model. We have shown that selecting multiple tokens based on their joint likelihood leads to better output perplexity. Thus, MTAD is more effective than multinomial sampling and vanilla speculative decoding. Furthermore, since MTAD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high. So at each iteration, MTAD can admit more draft tokens than vanilla speculative decoding, which results in better efficiency.
Next, we theoretically analyze the approximation error of MTAD. Lemma 3.4 shows the upper bound of MTAD’s perplexity. Theorem 3.5 reveals the upper bound of the ratio between the perplexity of approximate MTAD and exact MTJD. The proofs are given in Appendix A.
@@
$\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}})\geq$ (5)
$(1-\epsilon)\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\left(\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}})\right)$
@@
With the two assumptions above, when $N\rightarrow\infty$ we have
@@
With Theorem 3.6, we observe that when $q\rightarrow p$, we have $U\rightarrow 0$ and $\bar{\gamma}\rightarrow\infty$. Meanwhile, when the beam width for the auxiliary model is large enough, $\epsilon\rightarrow 0$, and the ratio bound in Theorem 3.5 converges to 1, This implies that MTAD converges to MTJD under these limiting conditions.
Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to produce multiple draft sequences and by applying tree-based attention (Miao et al., 2023) for simultaneous verification. However, our preliminary experiment results suggest that since MTAD already selects the longest accepted prefix sub-sequence, the advantage of generating multiple draft tokens is less significant. Moreover, this approach increases the memory cost during inference and may affect the error bounds derived above. Therefore, we leave a more detailed exploration of this extension as future work.